Determine the combined present value as of December 31, 2021, of the following four payments to be received at the end of each of the designated years, assuming an annual interest rate of 8%. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided. Round your final answers to nearest whole dollar amount.)

Year Received i = n = Payment Present Value
2022 8% 1 $9,500
2023 8% 2 10,050
2025 8% 4 11,600
2027 8% 6 13,950
Total $0
Answer #1
Year Received i= n= Payment Present Value
2022 8% 1 9500 8796 =9500*0.92593
2023 8% 2 10050 8616 =10050*0.85734
2025 8% 4 11600 8526 =11600*0.73503
2027 8% 6 13950 8791 =13950*0.63017
Total 34729